Solve the Cauchy Problem,

xux+(x+y)uy=u+1 with u(x,y)=x2 on y=0


1
Expert's answer
2021-12-15T07:57:26-0500

Taking

"x u_{x}+y u_{y}=0 \text { with } u(x, y)=x \text { on } x^{2}+y^{2}=1"

Due to the contour conditions we will consider the change of variables

"\begin{aligned}

&x=r \cos \theta \\

&y=r \sin \theta

\end{aligned}"

with

"\begin{aligned}

&d x=d r \cos \theta-r \sin \theta d \theta \\

&d y=d r \sin \theta+r \cos \theta d \theta

\end{aligned}"

we have

"\begin{aligned}

&u_{x}=u_{r} \frac{d r}{d x}+u_{\theta} \frac{d \theta}{d x} \\

&u_{y}=u_{r} \frac{d r}{d y}+u_{\theta} \frac{d \theta}{d y}

\end{aligned}"

Here from the characteristic curves

 

"\frac{d x}{x}=\frac{d y}{y} \Rightarrow \frac{d y}{d x}=\frac{y}{x}=\tan \theta"

So, we obtain

"x u_{x}+y u_{y}=0 \Longleftrightarrow 2 r u_{r}=0 \Rightarrow u(r, \theta)=C+\Phi(\theta)"

Note that

"\begin{aligned}

\frac{d r}{d x} &=\frac{1}{\cos \theta} \\

\frac{d r}{d y} &=\frac{1}{\sin \theta} \\

\frac{d \theta}{d x} &=0 \\

\frac{d \theta}{d y} &=0

\end{aligned}"

Now with the boundary conditions

"u(1, \theta)=\cos \theta \Rightarrow \Phi(\theta)=\cos \theta \text { and } C=0"

and finally

"u(r, \theta)=\cos \theta"

or in (x, y) coordinates, "u(x, y)=\frac{x}{\sqrt{x^{2}+y^{2}}}"